# Efficient Iterative Proximal Variational Inference Motion Planning

### v3

@@ 7.1 Efficiency Improvement through Parallel Computation
The first set of experiments is conducted to showcase the improved computation efficiency by leveraging the parallel computations on a GPU. Efficiency improvement is gained from parallel collision cost computation and the computing of marginal covariances using GBP.
(a) Parallel collision-checking on a dense trajectory. To demonstrate the improvement in collision-checking efficiency on a dense trajectory, we first compare the time of computing the expected collision factors between the proposed GPU parallel computation and a serial implementation [15]. We made the comparison on $2$D / $3$D point robots and $7$-DOF WAM arm robots with linear time-invariant (LTI) system
$A_{t}=\begin{bmatrix}0&I\\ 0&0\end{bmatrix},~{}a_{t}=\begin{bmatrix}0\\ 0\end{bmatrix},~{}B_{t}=\begin{bmatrix}0\\ I\end{bmatrix}.$ (43)
$2$D Point Robot $3$D Point Robot $7$-DoF WAM
@@
Figure 3: $7$-DOF WAM expected collision cost computation.
The results are shown in Tab. 1. On a dense trajectory (time discretization $N=750$), we achieved over $95\%$ improvement in the computation time compared with serial implementation on a CPU. Fig. 3 shows the computation time spent on the expected collision checking cost computation for different time discretizations. We observe that the parallel implementation is insensitive to the time discretization.
(b) GBP for computing the marginal covariances. Next, we demonstrate the efficiency improvement in computing the marginal covariances achieved through GBP. We compare with the brute-force way of computing the inverse of the precision matrix, $\Sigma_{\theta}^{-1}$, using a direct matrix inversion. The comparison results are in Tab. 2 and Fig. 4. We see a linear dependence on the number of time discretizations for GBP, and a polynomial ($N^{3}$) dependence of direct inversion. A $99.5\%$ improvement in implementation time is gained through GBP for a $7$-DOF WAM.
$2$D Point Robot $3$D Point Robot $7$-DoF WAM
@@
(c) The overall optimization time. Finally, we show the overall efficiency of the proposed paradigm by recording the total optimization time, and comparing with CPU serial implementation [14, 15]. The results are recorded in Tab. 3. We gained over $97\%$ improvement for the $7$-DOF planning problems.
$2$D Point Robot $3$D Point Robot $7$-DoF WAM
@@
Figure 7: Factorized and total costs in Fig. 5.
7.2 Planning Results for the 7-DOF Robot Arms
Next, we show the motion planning result for $2$ models of $7$-DOF robot arms, the WAM robot Arm [54], and the right arm of the PR2 robot [55]. We then visualize the results in the Robot Operating System (ROS) [56] with a high-fidelity robot model for both models integrated into the Moveit [57] motion planning package. Fig. 8 shows the planning results for the WAM robot arm, and Fig. 9 shows the results for the PR2 robot arm planning results. The experiment settings are listed in Tab. 7
7.3 Iterative PGVIMP for nonlinear dynamical system
We validate our proposed algorithm on a planar quadrotor system defined as
where $g$ is the gravity, $m$ represents the mass of the planar quadrotor, $l$ is the length of the body, and $J_{q}$ is the moment of inertia. $u_{1}$ and $u_{2}$ are the two thrust inputs to the system. In all our experiments, we set $m=1/\sqrt{2},l=\sqrt{2}$, and $J_{q}=1$. We linearize the system (44) around an initial nominal mean trajectory and perform the P-GVIMP algorithm on the obtained LTV system.
Iteration $1$ $3$ $10$ $20$
Norm difference $286.40$ $16.61$ $4.33$ $2.96$
Table 4: Norm of trajectory difference between two consecutive linearizations of the nonlinear dynamics.
Figure 8: Planning results for the WAM Robot for two tasks with $N=750$ states. The results are obtained from two algorithms: GPMP2 (blue) and KL-Proximal (Gray). For KL-Proximal, the settings are $T=2.0$ and $2.75$, $\Sigma_{obs}=20I$ and $19I$, The ratio between temperatures are $\hat{T}_{high}/\hat{T}_{low}=10$ and $15$, respectively, and $r+\epsilon_{sdf}=0.21$
@@ 7.3 Iterative PGVIMP for nonlinear dynamical system
(a) Convergence for LTV system. We first show the results for a linearized planar quadrotor system, using Algorithm 1. We show the convergence of the algorithm via a downsampled plot of the intermediate proposal Gaussian trajectory distributions in Fig. 5 with the low and high temperatures $\hat{T}$ in (12), and the corresponding cost evolutions, both for prior and collision factorized costs, and the total cost, in Fig. 7. After obtaining a collision-free trajectory with low temperatures, we switched to a high-temperature phase to put more weight on the entropy costs. The total costs keep decreasing throughout both processes.
WAMExp 1 WAMExp 2 PR2Exp 1 PR2Exp 2
P-GVIMP $10.26$s $9.48$s $14.03$s $11.86$s
GPMP2 $0.75$s $1.36$s $2.74$s $0.78$s
Table 5: The optimization time comparison for P-GVIMP and GPMP2 algorithms in the 7-DOF WAM and PR2 experiments with $n=750$ support states, averaged over 50 runs. For a 7-DOF robot, the GPMP2 algorithm has $14\!\times\!750=10,500$ optimization variables, while the KL-Proximal GVI has $14\!\times\!750+(3\!\times\!750\!-\!2)\times(14\!\times\!14)=443,608$ optimization variables.
@@
(b) Planning results for the nonlinear planar quadrotor. Next, we consider the full nonlinear planar quadrotor system (44) using the iterative-PGVIMP planning, i.e., Algorithm 2. We compute the norm of the trajectory difference between consecutive iterations, and the iterations are stopped once the norm difference is below a threshold. The results are recorded in Tab. 4, and the planning results are shown in Fig. 6.
(c) More experiments with different settings. We did experiments for $4$ different settings. We have $1$ obstacle in the environment. The start and goal states in the $4$ experiments are specified in the following Tab. 8.
@@
Fig. 11 shows the results of the nonlinear system of the planar quadrotor (44) for the $4$ experiment settings.
@@
(d) Go through or go around a narrow gap? Robust motion planning through entropy regularization. As motion planning with obstacles is a naturally non-convex problem, the solution is multi-modal. In the decision-making for motion planning, the risk and robustness of the solution are taken into consideration in our formulation by introducing the entropy of $q_{\theta}$ into the objective. This section shows an experiment for the planar quadrotor to fly through a narrow gap [14]. Two motion plans are obtained for the same motion planning task, as shown in Fig. 10. One plan (go-through plan) is visually riskier than the other (go-around plan). Our formulation provides a computable metric to compare the total cost consisting of optimality and robustness, between the two plans. In Tab. 6, we compute the sum of the prior and the collision costs, the entropy costs, and the total costs. By introducing lower entropy costs, our method selects the plan with lower risk over a shorter but riskier plan.
@@
Table 6: Comparing costs for two plans in Fig.10. The ‘MP’ is short for the sum of prior and collision costs, representing the optimality, and entropy represents the robustness.
(a) Experiment $1$.
(b) Experiment $2$.
(c) Experiment $3$.
(d) Experiment $4$.
Figure 11: Results of the i-PGVIMP for the nonlinear planar quadrotor dynamical system. $N=50$ support states are used, time span $T=4.0\sim 6.0$, radius $r+\epsilon_{sdf}=1.5$, $\Sigma_{obs}=6.0I\sim 7.5I$, low temperature $\hat{T}_{low}=1.0$, high temperature $\hat{T}_{high}=5.0$.
@@
In this work, we propose a distributed Gaussian Variational Inference approach to solve motion planning under uncertainties. The optimal trajectory distribution to a stochastic control problem serves as the target posterior in a variational inference paradigm. We leveraged this inference’s underlying sparse factor graph structure and proposed a distributed computation framework to solve the VI problem in parallel on GPU. Numerical experiments show the effectiveness of the proposed methods on an LTV system, and comparison studies demonstrated the computational efficiency.
@@ Appendix A Proof of Theorem 1
First take the gradient of the objectives in (17) with respect to $\mu_{\theta}$ and let it to be zero, we have
@@
from which we have the update for the mean as
@@
Since $\nabla_{\Sigma_{\theta}^{-1}}(\cdot)=-\Sigma_{\theta}\nabla_{\Sigma_{\theta}}(\cdot)\Sigma_{\theta}$, the necessary condition for the update rule for $\Sigma_{\theta}^{-1}$ can be derived by taking the gradients of the objectives in (17) with respect to $\Sigma_{\theta}$ and letting it to be zero, leading to
00 0 $=\nabla_{\Sigma_{\theta}}\psi_{d}(\theta_{k})+\frac{1}{2}(\frac{1}{\beta_{k}}\Sigma_{\theta_{k}}^{-1}+\mathbf{K}^{-1}-(1+\frac{1}{\beta_{k}})\Sigma_{\theta_{k+1}}^{-1}),$
from which we obtain the update rule for the $\Sigma_{\theta}^{-1}$ as
@@
Appendix B PGVIMP Algorithm Implementation Details
We present implementation details of the PGVIMP Algorithm 1.
(a) Step size selection. The proximal point algorithm is one type of gradient descent algorithm, which does not guarantee that the updated distribution at each step remains inside the distribution family $\mathcal{Q}$, especially with the step sizes being too large. For instance, the covariance updates in the Gaussian case does not guarantee that the updated covariances are always positive definite. We control the step size $\beta_{k}$ to enhance convergence efficiency by setting an upper bound on the KL divergence between the updated and current distributions. By solving the convex optimization problem
@@ Appendix B PGVIMP Algorithm Implementation Details
[…]
we are able to find the largest step size $\beta_{k}$ that satisfies the constraints, ensuring both stability and efficient convergence. In practice, we use a bi-section algorithm to choose a suitable stepsize that keeps the next step distribution close to the current one.
Appendix C Experiment Settings
We record the settings for different experiments we conducted in this work in Tab. 7 and Tab. 8.
Start & Goal
WAMExp 1 $[-0.8,-1.70,1.64,1.29,1.1,-0.106,2.2]$
$[0,0.94,0,1.6,0,-0.919,1.55]$
WAMExp 2 $[-0.9,-1.7,1.34,1.19,0.8,-0.126,2.5]$
$[-0.7,-1.35,1.2,1.0,-0.7,-0.1,1.2]$
PR2Exp 1 $[-0.7,0.175,-0.785,-1.05,-2.88,-0.6,0.785]$
$[-0.174,0.785,-0.26,-0.785,1.57,-0.35,1.57]$
PR2Exp 2 $[-1.9,0.15,-2.18,-0.524,3.14,-0.524,-1.047]$
$[0,0.171,0,-0.96,3.14,-0.785,0]$
Table 7: $7$-DOF Robot Arm Planning Experiment Settings.
Start Goal
[…]
Table 8: Planar Quadrotor Planning Experiment Settings.

### v4

@@ 7.1 Efficiency Improvement through Parallel Computation
The first experiment aims to demonstrate the improved computational efficiency achieved by parallel collision-checking on a GPU and marginal covariance computation using GBP.
(a) Parallel collision-checking. To demonstrate the improvement in collision-checking efficiency on dense trajectories, we compare the computation time of expected collision factors between the proposed GPU-parallel implementation and a serial baseline [16]. The experiments involve 2D and 3D point robots, as well as a 7-DOF WAM arm robot, modeled as a linear time-invariant (LTI) system
$A_{t}=\begin{bmatrix}0&I\\ 0&0\end{bmatrix},~a_{t}=\begin{bmatrix}0\\ 0\end{bmatrix},~B_{t}=\begin{bmatrix}0\\ I\end{bmatrix}$.
22D Point Robot 33D Point Robot 77-DoF WAM
@@
Figure 3: 77-DOF WAM expected collision cost computation.
The results appear in Table 2. On a dense trajectory with time discretization $N=750$, the GPU-parallel implementation achieves over $95\%$ reduction in computation time compared with the CPU-based serial implementation. Figure 3 reports the computation time for expected collision-checking costs under different time discretizations. The parallel implementation exhibits substantially lower sensitivity to the discretization level.
(b) GBP for computing the marginal covariances. Next, we demonstrate the efficiency gains in computing marginal covariances using GBP. Our proposed GBP approach is compared against the brute-force approach of directly inverting the precision matrix $\Sigma_{\theta}^{-1}$. The comparison results appear in Table 3 and Figure 4. GBP exhibits linear dependence on time discretizations $N$, whereas direct inversion scales cubically ($N^{3}$). For a 7-DOF WAM, GBP achieves a $99.5\%$ improvement in efficiency.
22D Point Robot 33D Point Robot 77-DoF WAM
@@
(c) The overall optimization time. Finally, the overall efficiency of the proposed paradigm is evaluated by recording the total optimization time and comparing it against a CPU-based serial implementation [15, 16]. The results appear in Table 4. For the 7-DOF planning problems, the proposed approach achieves over $97\%$ improvement in efficiency.
22D Point Robot 33D Point Robot 77-DoF WAM
@@
7.2 Planning Results for the 7-DOF Robot Arms
Next, the planning results for the WAM robot arm [52] and the right arm of the PR2 robot [53] are shown using the Robot Operating System (ROS) [54] and the Moveit [55] planning package. Figure 6 shows the planning results for the WAM robot arm, and Figure 5 shows the planning results for the PR2 robot arm.
The performance of P-GVIMP under uncertainty is further evaluated by comparing it with GPMP2 [12], BIT* [41], and FMT* [40]. To simulate perception noise, motion plans are first computed in a nominal obstacle environment. The obstacle positions are then perturbed by adding Gaussian noise, after which the plans are executed and the minimum distances from the trajectories to the obstacles are calculated. Figure 10 illustrates the trajectories generated by all four planners in the same perturbed environment, while Table 5 reports the average minimum distance from obstacles, computed over 50 randomly perturbed obstacle configurations.
Planner P-GVIMP GPMP2 BIT* FMT*
Average Distance 0.0216 $0.0052$ 0.0324-0.0324 0.0313-0.0313
Table 5: Average minimum distance from obstacles for different planners over 50 randomly perturbed environments.
7.3 Iterative P-GVIMP for nonlinear dynamical system
Our proposed algorithm is validated on a planar quadrotor
where $g$ is the gravity, $m$ represents the mass of the planar quadrotor, $l$ is the length of the body, and $J_{q}$ is the moment of inertia. $u_{1}$ and $u_{2}$ are the two thrust inputs to the system. In all experiments, $m=1/\sqrt{2}$, $l=\sqrt{2}$, and $J_{q}=1$. The system in (35) is linearized around a nominal trajectory, and the P-GVIMP algorithm is applied to the resulting LTV system.
Iteration 11 33 $10$ 2020
Norm difference $286.40$ $16.61$ $4.33$ $2.96$
Table 6: Norm difference between two consecutive linearizations of the nonlinear dynamic.
Figure 6: Planning results for the WAM Robot for two tasks with $N=750$ states. The results are obtained from two algorithms: GPMP2 (blue) and KL-Proximal (Gray). For KL-Proximal, we set $T=2.0$ and $2.75$, $\Sigma_{obs}=20I$ and $19I$, The ratio between temperatures are $\hat{T}_{high}/\hat{T}_{low}=10$ and $15$, respectively, and $r+\epsilon_{sdf}=0.21$
@@ 7.3 Iterative P-GVIMP for nonlinear dynamical system
Figure 9: Factorized and total costs in Figure 7.
(a) P-GVIMP
(b) GPMP2
(c) BIT*
(d) FMT*
Figure 10: PR2 planning results in an environment with perturbed obstacles. Each subfigure shows the resulting trajectory computed by the corresponding planner: (a) P-GVIMP, (b) GPMP2, (c) BIT*, and (d) FMT. Green obstacles represent the nominal environment, while yellow obstacles indicate the perturbed environment.
@@
(a) Empirical convergence for LTV system. We first present results for a linearized planar quadrotor system using Algorithm 1. Convergence is illustrated in Figure 7, which shows a down-sampled plot of the intermediate solution trajectory distributions with both low and high temperatures $\hat{T}$. The corresponding cost evolutions—including prior costs, collision-factorized costs, and total costs—are reported in Figure 9. After obtaining a collision-free trajectory under low temperatures, the algorithm transitions to a high-temperature phase to emphasize entropy costs. The total cost decreases monotonically throughout both phases.
@@
(b) Planning results for the nonlinear planar quadrotor. Next, consider the full nonlinear planar quadrotor system (35) using the iterative-P-GVIMP planning, i.e., Algorithm 2. The norm of the trajectory difference between consecutive iterations is computed, and the iterations are stopped once the norm difference is below a threshold. The results are recorded in Table 6, and the planning results are shown in Figure 8.
@@
(c) More experiments with different settings. We conducted experiments under four different settings with obstacles in the environment. Figure 12 shows the results of the nonlinear system of the planar quadrotor (35) for the $4$ experiment settings.
@@
(d) Go through or go around a narrow gap? Robust motion planning through entropy regularization. As motion planning with obstacles is a non-convex problem [56], and the solution is multi-modal. Risk and robustness of the solution are considered in our formulation by introducing the entropy of $q_{\theta}$ into the objective. This section shows an experiment for the planar quadrotor to fly through a narrow gap [15]. Two motion plan modes are obtained, as shown in Figure 11. One plan (go-through) is visually riskier than the other (go-around).
Our formulation provides a quantitative metric for comparing the optimality and robustness of the two plans through the entropy cost. Table 7 reports the sum of the prior and collision costs, the entropy costs, and the total costs. By incorporating lower entropy costs, the proposed method favors safer plans over shorter but riskier ones.
@@
Table 7: Comparing costs for two plans in Fig.11. The ‘MP’ is short for the sum of prior and collision costs, and entropy represents the robustness.
(a) Experiment 11.
(b) Experiment 22.
(c) Experiment 33.
(d) Experiment 44.
Figure 12: Results of the i-P-GVIMP for the nonlinear planar quadrotor dynamical system. $N=50$ support states are used, time span $T=4.0\sim 6.0$, radius $r+\epsilon_{sdf}=1.5$, $\Sigma_{obs}=6.0I\sim 7.5I$, low temperature $\hat{T}_{low}=1.0$, high temperature $\hat{T}_{high}=5.0$.
@@ 8 Conclusion
This work describes a distributed Gaussian Variational Inference approach to motion planning under uncertainties. The optimal trajectory distribution to a stochastic control problem serves as the target posterior in a variational inference paradigm. We leveraged this inference’s underlying sparse factor graph structure and proposed a distributed computation framework to solve the VI problem in parallel on GPU. Numerical experiments show the effectiveness of the proposed methods on an LTV system, and comparison studies demonstrated the computational efficiency.
@@ Appendix A Proof of Theorem 1
Taking the gradient of the objective in (14) with respect to $\mu_{\theta}$ and setting it to zero, we obtain
@@
which yields the update for the mean
@@
Since $\nabla_{\Sigma_{\theta}^{-1}}(\cdot)=-\Sigma_{\theta}\,\nabla_{\Sigma_{\theta}}(\cdot)\,\Sigma_{\theta}$, the necessary condition for updating $\Sigma_{\theta}^{-1}$ follows by computing the gradient of the objective in (14) with respect to $\Sigma_{\theta}$, setting it to zero, and obtaining
$0$ $=\nabla_{\Sigma_{\theta}}\psi_{d}(\theta_{k})+\frac{1}{2}(\frac{1}{\beta_{k}}\Sigma_{\theta_{k}}^{-1}+\mathbf{K}^{-1}-(1+\frac{1}{\beta_{k}})\Sigma_{\theta_{k+1}}^{-1}),$
from which the update rule for $\Sigma_{\theta}^{-1}$ follows
@@
Appendix B Theoretical Analysis of P-GVIMP
To establish the approximation accuracy and convergence guarantees of the proposed P-GVIMP algorithm, we define $\operatorname{KL}\left(\theta_{k+1}\!\parallel\!\theta_{k}\right):=\operatorname{KL}\left(q(x\!\mid\!\theta_{k+1})\!\parallel\!q\left(x\!\mid\!\theta_{k}\right)\right)$ and make the following assumptions:
(A1): The ELBO $\mathcal{L}$ is continuous and admits a maximum.
(A2): The function $\psi_{d}$ is L-smooth in $\mathcal{S}$
$\left\|\nabla\psi_{d}(\boldsymbol{\theta})-\nabla\psi_{d}\left(\boldsymbol{\theta}^{\prime}\right)\right\|\leq L\left\|\boldsymbol{\theta}-\boldsymbol{\theta}^{\prime}\right\|,\forall\boldsymbol{\theta},\boldsymbol{\theta}^{\prime}\in\mathcal{S}$
(A3): The function $\psi_{e}$ is convex.
(A4): There exists an $\alpha>0$ such that
$\left(\theta_{k+1}-\theta_{k}\right)^{T}\nabla\operatorname{KL}\left(\theta_{k+1}\!\parallel\!\theta_{k}\right)\!\geq\alpha\left\|\theta_{k+1}-\theta_{k}\right\|^{2}$
Approximation Error Bound
Proof.
Let $\delta:=\theta-\theta_{k}$, and define the first-order error
$\varepsilon(\theta)=\psi_{d}(\theta)-\left[\psi_{d}\left(\theta_{k}\right)+\nabla\psi_{d}\left(\theta_{k}\right)^{T}\delta\right].$
By the $L$-smoothness assumption on $\psi_{d}$ (A2),
$\psi_{d}\left(\theta\right)\leq\psi_{d}\left(\theta_{k}\right)+\nabla\psi_{d}\left(\theta_{k}\right)^{T}(\theta-\theta_{k})+\frac{L}{2}\left\|\theta-\theta_{k}\right\|^{2}$
Similarly, since $\psi_{d}$ is also $L$-smooth,
$\psi_{d}\left(\theta\right)\leq-\psi_{d}\left(\theta_{k}\right)-\nabla\psi_{d}\left(\theta_{k}\right)^{T}(\theta-\theta_{k})+\frac{L}{2}\left\|\theta-\theta_{k}\right\|^{2}$
Combining the two bounds gives
$|\varepsilon(\theta)|\leq\frac{L}{2}\|\delta\|^{2}\quad\Longrightarrow\quad\varepsilon(\theta)=\mathcal{O}\left(\|\delta\|^{2}\right).$
which shows that the first-order approximation error is of order $\mathcal{O}(\|\delta\|^{2})$ over any region where $\psi_{d}$ is $L$-smooth. ∎
Convergence Guarantee
Proof.
Given (A3) and (A4), Lemma 1 in [27] yields
$\left(\theta_{k}-\theta_{k+1}\right)^{T}\nabla\psi_{d}(\theta_{k})\geq\frac{\alpha}{\beta}\left\|\theta_{k+1}-\theta_{k}\right\|^{2}+\psi_{e}\left(\theta_{k+1}\right)-\psi_{e}\left(\theta_{k}\right).$
The $L$-smoothness assumption (A2) gives
$\psi_{d}\left(\theta_{k+1}\right)\leq\psi_{d}\left(\theta_{k}\right)+\nabla\psi_{d}\left(\theta_{k}\right)^{T}\left(\theta_{k+1}-\theta_{k}\right)+\frac{L}{2}\left\|\theta_{k+1}-\theta_{k}\right\|^{2}$
Combining the above two inequalities
$\psi_{d}(\theta_{k+1})+\psi_{e}(\theta_{k+1})$ $\leq\psi_{d}(\theta_{k})+\psi_{e}(\theta_{k})-\left(\frac{\alpha}{\beta}-\frac{L}{2}\right)\|\theta_{k+1}-\theta_{k}\|^{2}$
$\Rightarrow\quad\mathcal{L}(\theta_{k+1})$ $\leq\mathcal{L}(\theta_{k})-\left(\frac{\alpha}{\beta}-\frac{L}{2}\right)\|\theta_{k+1}-\theta_{k}\|^{2}.$
Therefore, for any step size $0<\beta<\frac{2\alpha}{L}$, the sequence $\{\mathcal{L}(\theta_{k})\}$ is strictly increasing and bounded by its maximum. Consequently, $\left(\frac{\alpha}{\beta}-\frac{L}{2}\right)\left\|\theta_{k+1}-\theta_{k}\right\|^{2}\to 0$, and hence $\left\|\theta_{k+1}-\theta_{k}\right\|\to 0$. ∎
Appendix C Implementation Details
This section presents the implementation details of the P-GVIMP Algorithm 1.
Step size selection
The proximal point algorithm does not guarantee that the updated distribution at each step remains inside the distribution family $\mathcal{Q}$, especially when the step sizes $\beta_{k}$ are too large. The updates do not guarantee that the updated covariances are always positive definite. To enhance convergence efficiency, we control $\beta_{k}$ by imposing an upper bound on the KL divergence between the updated and current distributions. To this end, we solve the convex program
@@ Appendix C Implementation Details
$\text{s.t.}\quad{\rm KL}\left(q_{\theta_{k+1}}(\mathbf{x})\|q_{\theta_{k}}(\mathbf{x})\right)\leq\epsilon,\;\Sigma^{-1}_{\theta_{k+1}}\succ 0.$
In practice, a bi-section algorithm is used to choose a step size that keeps the next step distribution close to the current one.